I'm struggling with a homework question.
Its reads as follows
Write an equation for each of the following.
a) The conic with center at the origin, focus a (0,-3) and e= 3/5
b) The conic centered at the origin with vertex at (2,0) and e=3/2

Answer by stanbon(48581)

Write an equation for each of the following.
a) The conic with center at the origin, focus a (0,-3) and e= 3/5
Plot the center and the focus point.
Since e < 1 the conic is an ellipse.
Since the focus is below the center the form is
x^2/b^2 + y^2/a^2 = 1
-----------------------------
Since the distance from the center to the focus is 3, c=3.
-----
Since e = c/a = 3/5, a = 5
--------------------------
Since a^2 = b^2+c^2, 25 = b^2 + 9
b = sqrt(16) = 4
---------------------
Equation:
x^2/16 + y^2/25 = 1
============================
b) The conic centered at the origin with vertex at (2,0) and e=3/2
------
Plot the center and the vertex.
--------
Since e > 1 the conic is a hyperbola.
-------------
Since the focus is to the right of the center,
the hyperbola opens to the right and to the left.
-------------
Form: x^2/a^2 - y^2/b^2 = 1
-------------------------------
Distance from the center to the focus = c = 2.
-----------
Since e = c/a = 3/2, a = 4/3
------------
Since c^2 = a^2+b^2, 4 = (4/3)^2 + b^2
b^2 = 4 - (16/9) = 20/9
-----------------------
Equation:
x^2/(16/9) - y^2/(20/9) = 1
9x^2/16 - 9y^2/20 = 1
============================
Cheers,
Stan H.